# Function-Correcting Codes for Insertion-Deletion Channels

### v1

@@ VII Locally $(\lambda,\rho)_{I}$-function
Let $f$ be a locally $(\lambda,2t)_{ID}$ bounded function. Then by Lemma 19, there exists a mapping
@@
$\mathrm{Enc}:\mathbb{F}_{2}^{k}\to\mathbb{F}_{2}^{k+N_{ID}^{(2)}(\lambda,2(t+k);k)}$
@@
$\mathrm{Enc}(u)=(u,u_{p}),\quad\text{where }u_{p}=C_{\mathrm{Col}_{f}(u)}.$
@@
Hence, in both cases, the encoding satisfies the desired distance property, proving that $\mathrm{Enc}$ defines an FCIDC with redundancy $r=N_{ID}^{(2)}((\lambda,2t);k)$. ∎
Corollary 5.
@@
The lower bound follows from Corollary 1, while the upper bound is obtained from Theorem 3 together with the observation that $N^{(2)}_{ID}(2,2(t+k);k)=t+k$. ∎
Corollary 6.
@@
From Lemma 19, there exists a mapping $\mathrm{Col}_{f}:\mathbb{F}_{2}^{k}\rightarrow[3]$ corresponding to the function $f$, such that for any $u,v\in\mathbb{F}_{2}^{k}$,
$\mathrm{Col}_{f}(u)\neq\mathrm{Col}_{f}(v)\quad\text{whenever}\quad f(u)\neq f(v)\ \text{and}\ d_{I}(u,v)\leq 2t.$
@@
$Enc:\mathbb{F}_{2}^{k}\rightarrow\mathbb{F}_{2}^{k+2(k+t)}$
defined as follows: $Enc(u)=(u,u_{p})$ where $u_{p}=(u_{p}^{\prime})^{t+k}$ and
@@
It can easily be seen that $LCS((00)^{t+k},(01)^{t+k})=t+k$ as both the vectors have $t+k$ common $0s$. Similarly, one can easily see $LCS((11)^{t+k},(01)^{t+k})=t+k$ and $LCS((00)^{t+k},(11)^{t+k})=0$. Therefore, the insdel distance between any distinct redundancy vector is greater than or equal to $2(t+k)$. We now prove that the above encoding function $Enc$ is an FCIDC with redundancy $2(t+k)$.
@@
Then we can find two vectors $y_{1},y_{2}\in\mathbb{F}_{2}^{k}$ with $d(x,y_{1})\leq 2t$, $d(x,y_{2})\leq 2t$, and
@@
$f(y_{1})-f(y_{2})\geq 4t+1.$ (2)
On the other hand, using the results of Lemma 2 and the triangle inequality, we get,
${r}(y_{1})-r(y_{2})\leq d_{ID}(y_{1},y_{2})\leq d_{ID}(x,y_{1})+d(x,y_{2})\leq 4t.$ (3)
Thus from equation 2 and 3 we have,
@@
VIII Conclusion

### v2

@@ 7 Locally $(\lambda,\rho)_{ID}$-function
Let $f$ be a locally $(\lambda,2t)_{ID}$-function. Then by Lemma 7.3, there exists a mapping
@@
$\psi:\mathbb{F}_{2}^{k}\to\mathbb{F}_{2}^{k+N_{ID}^{(2)}(\lambda,2(t+k);k)}$
@@
$\mathrm{\psi}(u)=(u,u_{p}),\quad\text{where }u_{p}=C_{\mathrm{Col}_{f}(u)}.$
@@
Hence, in both cases, the encoding satisfies the desired distance property, proving that $\psi$ defines an FCIDC with redundancy $r=N_{ID}^{(2)}((\lambda,2t);k)$. ∎
Corollary 7.5.
@@
The lower bound follows from Corollary 3.19, while the upper bound is obtained from Theorem 7.4 together with the observation that $N^{(2)}_{ID}(2,2(t+k);k)=t+k$. ∎
Corollary 7.6.
@@
From Lemma 7.3, there exists a mapping $\mathrm{Col}_{f}:\mathbb{F}_{2}^{k}\rightarrow[3]$ corresponding to the function $f$, such that for any $u,v\in\mathbb{F}_{2}^{k}$,
$\mathrm{Col}_{f}(u)\neq\mathrm{Col}_{f}(v)\quad\text{whenever}\quad f(u)\neq f(v)\ \text{and}\ d_{ID}(u,v)\leq 2t.$
@@
$\psi:\mathbb{F}_{2}^{k}\rightarrow\mathbb{F}_{2}^{k+2(k+t)}$
defined as follows: $\psi(u)=(u,u_{p})$ where $u_{p}=(u_{p}^{\prime})^{t+k}$ and
@@
It can easily verified that $LCS((00)^{t+k},(01)^{t+k})=t+k$ as both the vectors have $t+k$ common $0s$. Similarly, $LCS((11)^{t+k},(01)^{t+k})=t+k$ and $LCS((00)^{t+k},(11)^{t+k})=0$. Therefore, the insdel distance between any distinct redundancy vector is at least $2(t+k)$. We now prove that the above encoding function $\psi$ is an FCIDC with redundancy $2(t+k)$.
@@
Then we can find two vectors $y_{1},y_{2}\in\mathbb{F}_{2}^{k}$ with $d_{ID}(x,y_{1})\leq 2t$, $d_{ID}(x,y_{2})\leq 2t$, and
@@
$f(y_{1})-f(y_{2})\geq 4t+1.$ (7.1)
On the other hand, using the results of Lemma 2.9 and the triangle inequality, we get,
${r}(y_{1})-r(y_{2})\leq d_{ID}(y_{1},y_{2})\leq d_{ID}(x,y_{1})+d_{ID}(x,y_{2})\leq 4t.$ (7.2)
Thus from equation 7.1 and 7.2 we have,
@@
Function Lower Bound Upper Bound
VT-syndrome function $\frac{2t(k+1)}{k+2}$ $\Bigg\lceil\frac{\ln(k+1)+2(t+k-1)\ln\Big(\frac{ek}{t+k-1}\Big)}{\ln 2}\Bigg\rceil$.
Number of runs function $\frac{2}{(t+2)^{2}}\left[\frac{5t(t+1)(t+2)}{6}+\left\lceil\frac{t}{2}\right\rceil(t+1)+\left\lceil\frac{t}{2}\right\rceil^{2}\right]$ $\Bigg\lceil\frac{\ln(2t+1)+2(t+k-1)\ln\Big(\frac{ek}{t+k-1}\Big)}{\ln 2}\Bigg\rceil$
Maximum run-length function $\frac{2}{(t+2)^{2}}\left[\frac{5t(t+1)(t+2)}{6}+\left\lceil\frac{t}{2}\right\rceil(t+1)+\left\lceil\frac{t}{2}\right\rceil^{2}\right]$ $\Bigg\lceil\frac{\ln(k)+2(t+k-1)\ln\Big(\frac{ek}{t+k-1}\Big)}{\ln 2}\Bigg\rceil$
Locally binary insdel function $t$ $t+k$
Table 1: Lower bound and upper bound on the optimal redundancy of FCIDCs for different functions.
8 Conclusion
@@ Acknowledgment
9 Appendix
We show that the construction 1 yields an FCIDC for the redundancy vectors.
For $t=1$, assume $r(x_{i})\neq r(x_{j})$ and consider the following two cases.
Case 1: $r(x_{i})\equiv r(x_{j})\text{ }smod{3}$.
Since the run counts belong to the same residue class modulo 3 and are distinct, the smallest possible run count difference is 3.
From Lemma 2.9 we obtain
$d_{ID}\bigl(\psi(x_{i}),\psi(x_{j})\bigr)=d_{ID}\bigl((x_{i},p_{r(x_{i})}),(x_{j},p_{r(x_{j})})\bigr)\geq 2\Bigl\lceil\frac{3}{2}\Bigr\rceil=4.$
Thus, whenever $r(x_{i})\equiv r(x_{j})\text{ }smod{3}$ and $r(x_{i})\neq r(x_{j})$,
$d_{ID}\bigl(\psi(x_{i}),\psi(x_{j})\bigr)\geq 4.$
Case 2: $r(x_{i})\not\equiv r(x_{j})\text{ }smod{3}$.
Here the redundancy parts $p_{r(x_{i})}$ and $p_{r(x_{j})}$ are distinct codewords code with minimum insdel distance $2(k+1)$. Hence we have,
$d_{ID}\bigl(\psi(x_{i}),\psi(x_{j})\bigr)$ $=d_{ID}\bigl((x_{i},p_{r(x_{i})}),(x_{j},p_{r(x_{j})})\bigr)$
$\geq d_{ID}(x_{i},x_{j})+d_{ID}(p_{r(x_{i})},p_{r(x_{j})})-2\min\{k,2(k+1)\},$
Since, $d_{ID}(x_{i},x_{j})\geq 2$, $d_{ID}(p_{r(x_{i})},p_{r(x_{j})})\geq 2(k+1)$ and $\min\{k,2(k+1)\}=k$, we obtain
$d_{ID}\bigl(\psi(x_{i}),\psi(x_{j})\bigr)\geq 2+2(k+1)-2k=4.$
For $t=2$ a completely analogous analysis shows that the concatenation of message vectors with the corresponding redundancy vectors yields a FCIDC.
For $t\geq 3$, assume $r(x_{i})\neq r(x_{j})$ and consider the following two cases.
Case 1: $r(x_{i})\equiv r(x_{j})\text{ }smod{(2t+1)}$.
Since the run counts belong to the same residue class smodsmod $2t+1$ and are distinct, the smallest possible run count difference is $2t+1$.
From Lemma 2.9 we obtain
$d_{ID}\bigl(\psi(x_{i}),\psi(x_{j})\bigr)=d_{ID}\bigl((x_{i},p_{r(x_{i})}),(x_{j},p_{r(x_{j})})\bigr)\geq 2\Bigl\lceil\frac{2t+1}{2}\Bigr\rceil=2t+2.$
Thus, whenever $r(x_{i})\equiv r(x_{j})\text{ }smod{(2t+1)}$ and $r(x_{i})\neq r(x_{j})$,
$d_{ID}\bigl(\psi(x_{i}),\psi(x_{j})\bigr)\geq 2t+2.$
Case 2: $r(x_{i})\not\equiv r(x_{j})\text{ }smod{(2t+1)}$.
Here the redundancy parts $p_{r(x_{i})}$ and $p_{r(x_{j})}$ are distinct codewords code with minimum insdel distance $2(k+t)$. Hence we have,
$d_{ID}\bigl(\psi(x_{i}),\psi(x_{j})\bigr)$ $=d_{ID}\bigl((x_{i},p_{r(x_{i})}),(x_{j},p_{r(x_{j})})\bigr)$
$\geq d_{ID}(x_{i},x_{j})+d_{ID}(p_{r(x_{i})},p_{r(x_{j})})-2\min\{k,2(k+1)\},$
Since, $d_{ID}(x_{i},x_{j})\geq 2$, $d_{ID}(p_{r(x_{i})},p_{r(x_{j})})\geq 2(k+1)$ and $\min\{k,2(k+1)\}=k$, we obtain
$d_{ID}\bigl(\psi(x_{i}),\psi(x_{j})\bigr)\geq 2+2(k+t)-2k=2t+2.$